OpenStudy (anonymous):

solve \[2^{6x-1}=28\] and check

OpenStudy (cruffo):

to start, take the natural log of both sides.

OpenStudy (anonymous):

so ... \[\ln 2^{6x-1} = \ln 28\] ?

OpenStudy (cruffo):

cool. Now use power rule to bring out the (6x-1)

OpenStudy (anonymous):

(6x-1)ln2=ln28

OpenStudy (cruffo):

excellent! now divide both sides by ln2

OpenStudy (anonymous):

uhh... 6x-1=4.807355

OpenStudy (cruffo):

you can leave this part as the exact answer - usually don't need to go to decimal answer

OpenStudy (cruffo):

6x-1 = ln(28)/ln(2)

OpenStudy (anonymous):

so leave it as ln28/ln2 ? is that what you mean?

OpenStudy (cruffo):

yep. It will depend on what format your answer need to be in. Do the instructions say to answer in terms of natural logs, or as a decimal?

OpenStudy (anonymous):

it just says to solve the equation for x and check the solution. the answer, however, is kept in terms of log.

OpenStudy (cruffo):

ok. So you should leave this part as ln28/ln2

OpenStudy (anonymous):

ok, i did up to that part, but now idk what to do... i have the answer, idk how to get to it.

OpenStudy (cruffo):

To finish solving for x, add one to both sides and divide both sides by 6 6x = ln28/ln2 + 1 x = ln28/(6ln2) + 1/6

OpenStudy (cruffo):

there are several ways to write this answer...

OpenStudy (anonymous):

my worksheet tells me that x=log56/6log2

OpenStudy (cruffo):

Your answer and their answer are the same amount - check using a calc we started using ln but the solution is the same as using log. So ln28/(6ln2) + 1/6 is the same as log28/(6log2) + 1/6

OpenStudy (anonymous):

wait before we go any further, would it be ln28/(6ln2) or ln28/ln2 +1/6 ?? sorry i got lost

OpenStudy (cruffo):

both terms have to be divided by 6 \[\large 6x = \frac{\ln28}{\ln2} + 1\] \[\large \frac{1}{6} \cdot 6x = \frac{1}{6} \cdot \frac{\ln28}{\ln2} + \frac{1}{6} \cdot1\] \[\large x = \frac{\ln28}{6\ln2} + \frac{1}{6}\]

OpenStudy (anonymous):

oooh okay,

OpenStudy (anonymous):

how come it is log56/6log2 then? where did the log56 come from? :S

OpenStudy (cruffo):

To get this answer to look like their answer we just need to multiply the 1/6 by ln2/ln2 \[\large x = \frac{\ln28}{6\ln2} + \frac{1}{6}\cdot\frac{\ln2}{\ln2}\] \[\large x = \frac{\ln28}{6\ln2} + \frac{\ln2}{6\ln2}\] since the fractions have a common denominator, we can add \[\large x = \frac{\ln28 + ln2}{6\ln2} \] then use product rule to write the numerator as one ln

OpenStudy (anonymous):

oooh ! okay. that makes sense. sorry another question. when i plug in ln28/(6ln2)+1/6 and ln28/(6ln2)+1/6 i get .55162 and .23927 ... :S am i doing it wrong?

OpenStudy (cruffo):

???

OpenStudy (anonymous):

oh... LOL oooppsss. i put the eqn in wrong in my calculator then.. ._.

OpenStudy (anonymous):

k thanks @cruffo (:

OpenStudy (cruffo):

np